A chemist has one solution of 20% acid and another of 50% acid. How many milliliters of each are needed to obtain 18 milliliters of 30% acid?
Answer by richwmiller(17219)

x+y=18
.2x+.5y=.3*18
x=12 ml of 20%
y=6 ml of 50%
check
.2*12+.5*6=.3*18
2.4+3=5.4
5.4=5.4
ok